Circular No. 6351

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996N IN NGC 1398 A. Williams and R. Martin, Perth Observatory, report their discovery,

during the Perth Astronomy Research Group’s automated supernova search on the Perth-Lowell 0.61-m reflector, of a supernova (R ∼ 16) on Mar. 12.538 UT. SN 1996N is located 46′′ east and 12′′ north of the center of NGC 1398 (α = 3^{h}38^{m}.9, δ = −26^{o}20′, equinox 2000.0). Nothing was visible at this location on Feb. 16.

L. Germany and B. Schmidt, Mount Stromlo and Siding Spring Observatories; and R. Stathakis and H. Johnston, Anglo-Australian Observatory, report that a spectrogram (range 550–900 nm) obtained on Mar. 23.4 UT at the Anglo-Australian Telescope shows SN 1996N to be a type-Ib/c supernova, ∼ 2 weeks past maximum.

SUPERNOVA 1996B IN NGC 4357 L. Wang and J. C. Wheeler, University of Texas at Austin, report:

“SN 1996B was observed again using the Imaging Grism Instrument at the 2.1 m telescope of the McDonald Observatory on Mar. 20 UT. Inspection of a preliminarily-reduced CCD spectrum (range 424.7–813.5 nm, resolution 2.0 nm) shows that the Hα feature that was strong in our Jan. 22 spectrum (IAUC 6292) has now decreased in strength sharply. The residual Hα feature is flat-topped (with FWHM ∼ 10 000 km/s) and resembles that of SN 1993J in its late nebular phase. A weak absorption feature is still detectable at the blue side of the flat-topped Hα emission feature. There are also indications of the existence of the various helium lines that were detected in SN 1993J, but the spectrum is too noisy to establish this firmly. The redshift of the host galaxy, as measured from the neighboring H II region, is 4440 km/s. SN 1996B is therefore likely to be another type-IIb supernova similar to SN 1993J. Observations in other wavelengths are encouraged.”

COMET 29P/SCHWASSMANN-WACHMANN 1 Further total visual magnitude estimates (cf. IAUC 6320): Feb. 24.39

UT, 11.2 (C. S. Morris, Lockwood Valley, CA, 0.26-m reflector); Mar. 8.85, 11.1 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector); 13.29, 10.8 (R. Keen, Mt. Thorodin, CO, 0.32-m reflector); 17.25, 10.9 (Morris); 23.19, 11.3 (A. Hale, Cloudcroft, NM, 0.41-m reflector).

1996 March 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green